OpenStudy (anonymous):

Could someone plezzzzz help me! ill give a medal ! Add and then reduce to lowest terms. 7 3/5 + 4 1/10

OpenStudy (anonymous):

\[7\frac35 + 4\frac1{10} = \frac15(7*3 + \frac42) = \frac{23}5\]

OpenStudy (anonymous):

thank you so much :) !!!!

OpenStudy (anonymous):

no problem!